A rectangular piece of cloth with a length of 1.50 meters and a width of 0.80 meters must be divided into ten, equal-sized units for the project. What are the dimensions of the ten  smaller portions?


1
Expert's answer
2021-03-29T09:02:22-0400

Length:


"a=\frac{A}{10}=\frac{1.5}{10}=0.15\ m"

Width:


"b=\frac{B}{10}=\frac{0.8}{10}=0.08\ m"